For a city tour, Yolanda decides to order T-shirts for all the participants. It will cost $2 per shirt for the small size, $3 per shirt for the medium size. Yolanda ordered a total of 50 T-shirts and spend a total of $130. How many are small sized shirts? How many are medium sized shirts?
Define your variables with proper ‘let’ statements.
Write two linear equations that model the problem.
How many small size shirts?
How many medium size shirts?

Answer by Boreal(15235)

S+M=50
2S+3M=130, money equation
-2S-2M=-100
M=30 medium or $90
S=20 small or $40